ind all solutions to the equation on the interval (0,2) - 2 cos sinx+1=0
Answer #1

Given - 2cos(x) + sin(x) +1 = 0 To find all solutions of the equation on the interval o to 27 = > - 2cos”(x) + sin(x) + 1 = 0= > 2y(y + 1) – 1(y + 1) = 0 = > (2y – 1)(y + 1) = 0 = > y = , y = -1 = > sin(x) = } = > x = sin -11 = > x = 30° = , 150° = 5